OpenStudy (anonymous):

show that A(-3,1)B(1,2)C(0,-1)D(-4,-2)are vertices of a parellelogram

OpenStudy (anonymous):

show that diagonals AC and BD bisect each other.... u can show it by finding midpoint of AC and BD........ try using midpoint formula you will get the same points....

OpenStudy (anonymous):

A-3/2 ,C 1/2 B -1, D 1 IS THIS THE CORRECT ANSWER

OpenStudy (anonymous):

nope, how is that all numbers...........you need to find co-ordinates.....let me show u whole process

OpenStudy (anonymous):

midpoint of AC = ((-3+0)/2,(1-1)/2)=(-3/2,0) similarly midpoint of BD = ((1-4)/2,(2-2)/2)=(-3/2,0) Since the midpoints of diagonals coincide, the given vertices are of parallelogram....

OpenStudy (anonymous):

you can use second method by finding slope of all sides, in this case, slope of AB=CD and slope of BC= AD and formula for slope = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Hope you got it!

OpenStudy (anonymous):

THNNXXXX AM TRYING IT KNW